1.an oxide of iron consists of 70%by mass of iron and the rest is oxygen. Find i
ts simplest formula.
2.a gas contains 16.5g of nitrogen and 3.5g of hydrogen.what is it's simplest formula
4.calc the moles of electrons involved when :28g of iron are converted to iron (two)ions
11.5g of sodium are converted to sodium ions
1
Expert's answer
2020-04-02T08:59:41-0400

1. FexOy

w(Fe) = 70%; w(O) = 100% - 70% = 30%

m(FexOy) = 100 g

m(Fe) = 70 g; m(O) = 30 g

x : y = n(Fe) : n(O) = m(Fe)/M(Fe) : m(O)/M(O) = 70/56 : 30/16 = 1.25 : 1.875 = 2 : 3

Fe2O3

2. NxHy

x : y = n(N) : n(H) = m(N)/M(N) : m(H)/M(H) = 16.5/14 : 3.5/1 = 1.17857143 : 3.5 = 1 : 3

NH3

4. Fe = Fe2+ + 2e

n(Fe) = m(Fe)/M(Fe) = 28/56 = 0.5 mol

0.5 × 2 = 1.0 moles of electrons

Na = Na+ + e

n(Na) = m(Na)/M(Na) = 11.5/23 = 0.5 mol

0.5 × 1 = 0.5 moles of electrons